OpenStudy (anonymous):

Solve the following system of equations. 9x + 5y = -154 4x + y = -22

jimthompson5910 (jim_thompson5910):

the system is 9x + 5y = -154 4x + y = -22 right?

jimthompson5910 (jim_thompson5910):

or no?

jimthompson5910 (jim_thompson5910):

type here @jazzykaila33

jimthompson5910 (jim_thompson5910):

where it says "type your reply"

OpenStudy (anonymous):

okay sorry

OpenStudy (anonymous):

Joanna went school supply shopping. She spent $27.72 on notebooks and pencils. Notebooks cost $1.96 each and pencils cost $1.14 each. She bought a total of 20 notebooks and pencils. How many of each did she buy? 4 notebooks; 16 pencils 6 notebooks; 14 pencils 9 notebooks; 11 pencils 11 notebooks; 9 pencils

jimthompson5910 (jim_thompson5910):

that's fine, but what about the first one?

OpenStudy (anonymous):

the question got deleted so I cant pull it back up, sorry for the difficulties

jimthompson5910 (jim_thompson5910):

oh, I see, that's fine

jimthompson5910 (jim_thompson5910):

let n = # of notebooks p = # of pencils

jimthompson5910 (jim_thompson5910):

"She bought a total of 20 notebooks and pencils." so which equation can we form?

jimthompson5910 (jim_thompson5910):

sorry the site glitched

jimthompson5910 (jim_thompson5910):

I'm back

jimthompson5910 (jim_thompson5910):

let n = # of notebooks p = # of pencils "She bought a total of 20 notebooks and pencils." so which equation can we form?

jimthompson5910 (jim_thompson5910):

see here @jazzykaila33

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

ready

OpenStudy (anonymous):

not so sure but i think its standard form we can use Ax+by=c

jimthompson5910 (jim_thompson5910):

we have n notebooks and p pencils we have a total of 20 notebooks and pencils so what can we do with n and p to get 20?

OpenStudy (anonymous):

add?

jimthompson5910 (jim_thompson5910):

yes, add n and p to get 20 n+p = 20

jimthompson5910 (jim_thompson5910):

which variable do you want to isolate? n? or p?

OpenStudy (anonymous):

um. lets do n

jimthompson5910 (jim_thompson5910):

ok so we have n+p = 20 to isolate n, we need to move that p over

jimthompson5910 (jim_thompson5910):

how do we do this?

OpenStudy (anonymous):

make the numbers x out by subtracting its opposite?

jimthompson5910 (jim_thompson5910):

well just subtracting

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):